OpenStudy (anonymous):

Veronica has a lawn, ABCD. She has placed a watering hose, BD, as shown below. Part A: Veronica plans to put a fence along the length AD of her lawn. What is the length of the fence required? Part B: Using complete sentences, explain how you arrived at the answer for Part A.

OpenStudy (anonymous):

@Nali

OpenStudy (nali):

ok ur problem has two parts to solve part A

OpenStudy (nali):

first find the lenght of BD using the pathorgrean formula do u know it ?

OpenStudy (anonymous):

yes a^2+b^2=c^2

OpenStudy (nali):

ok so find BD and tell me what u get ?

OpenStudy (anonymous):

c^2=1225

OpenStudy (nali):

ok take the squarerrot of it to fin BD

OpenStudy (anonymous):

35

OpenStudy (nali):

ok

OpenStudy (nali):

do the same thing as before what would se to solve for AD sin or cos

OpenStudy (nali):

and remember what would they equal like opp/hyp

OpenStudy (anonymous):

sin?

OpenStudy (nali):

did u learn about tan yet?

OpenStudy (anonymous):

yes thats opp/adj

OpenStudy (nali):

ok good so in this case since we do not know what the hypostenuse is as a result, we use tan

OpenStudy (nali):

so how would set it up?

OpenStudy (nali):

tan 60=35/x and then just solve for x